# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

### v1

@@ III-B Multi-view Adversarial Learning
Objects of the same category share many similar structures, and therefore, their shape distributions are close. To make the generated complete point cloud as realistic and reasonable as possible, we develop an adversarial learning module to learn the category-specific shape knowledge from the real depth maps. Specifically, we first employ a silhouette renderer to render $\mathbf{P}_{out}$ and $\mathbf{P}_{c}$ to two silhouette maps $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ from the same viewpoint as $\mathbf{D}_{0}$. $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ are supervised by $\mathbf{D}_{0}$. Then we leverage a depth renderer to render $\mathbf{P}_{out}$ to depth maps from different viewpoints. These depth maps and those from the category-specific depth image bank are fed to a discriminator, which consists of five 2D Convolutional layers, to discriminate whether they are generated or real samples. It is important to note that the depth image bank is composed of single-view maps, each corresponding one-to-one with the input partial shape, without including samples from additional viewpoints.
@@
Our differentiable rendering pipeline is derived from [68]. Given the intrinsic and extrinsic camera parameters, [68] first transform the point cloud from world to camera coordinates and then project it into the 2D image plane. During rasterization, each point is rendered as a circular splat with a predefined radius in screen space, where the opacity of the splat decreases away from its center. The value of each pixel in the silhouette maps is computed by blending the information from the K nearest points along the z-axis whose splat regions overlap that pixel. For depth maps, each pixel’s value is the z-coordinate of the point that has the smallest z-value among the K nearest points.
@@
However, [68] assigns a fixed and identical radius to each point splat. In fact, the density of 2D points projected by a point cloud with a fixed point number from different viewpoints is different. For example, as shown in Figure 3, when there is a large foreground area of a rendered image, the density of the projected points is small (we call these low-density depth maps). In such cases, a small radius cannot cover all areas of the visible surface. Many “hole pixels” will appear in the rendered depth map. These “hole pixels” are inaccurately represented by the depth values of the points that should have been occluded. In contrast, in scenarios with few foreground pixels, the density of points is large (we call these high-density depth maps). At this time, a small radius suffices to cover the entire visible surface, and thus, the rendered depth map is more realistic and smooth. Therefore, a fixed point radius is inadequate for rendering under different viewpoints.
Intuitively, using a larger point radius for rendering might produce smoother depth maps in both high point density and low point density scenarios. However, in this way, the projected points cover a larger area on the 2D plane, potentially extending the edges of the foreground pixels beyond the original boundaries of the object. This results in a significant overestimation of the object’s size. Moreover, in high-density regions, a large point radius might cause multiple points to overlap excessively. This overlap could obscure finer details, as the rendered depth value for a pixel might be dominated by fewer points, causing the depth map to lose important structural information. Besides, an excessively large point radius may limit the rendering process to updating gradients for only a small subset of points. Therefore, the point radius should dynamically vary with the density of the projected 2D points. The higher the density, the smaller the radius, and vice versa.
@@
To this end, we introduce a density-aware radius estimation (DARE) algorithm to adaptively determine a suitable radius for points from each viewpoint. First, we predefine an initial point radius for the generated points and then render them into a coarse depth map. Note that the predefined point radius should cover the projected area. We then calculate the number of foreground pixels on this depth map, denoted as $A$, and subsequently, the density of the projected points on the depth map is defined as
@@
where $r$ denotes the point radius. Note that the parameter $\eta$ can be pre-estimated based on the initial point radii. Specifically, for each category, the density of projected points could be approximated using the average foreground pixel number of all the depth maps in the dataset, as specified in Formula 6. Following this, $\eta$ can be computed according to Formula 7. In this way, the radius of the point cloud dynamically adjusts in response to variations in the density of projected points from various viewpoints, resulting in a smoother rendered depth map.
@@ III-C Loss Functions
Furthermore, we use a rendering loss to supervise the predicted point cloud in 2D space. First, we binarize the depth image $\mathbf{D}_{0}$ to obtain the 2D silhouette ground truth $\mathbf{S}_{0}$ from the input viewpoint. Subsequently, we compute the Mean Squared Error (MSE) loss between $\mathbf{S}_{0}$ and $\mathbf{S}_{out}$. Note that due to the sparsity of the coarse point cloud $\mathbf{P}_{c}$, the foreground pixels of $\mathbf{S}_{c}$ are also sparse and discrete. In contrast, the foreground of $\mathbf{S}_{0}$ is continuous and dense. Therefore, we propose a Masked MSE loss, which only calculates the MSE loss on the foreground pixels of the rendered image. This trick confines the projected points of the rendered image to be accurately positioned.
@@
IV Experiments
IV-A Dataset and Evaluation Metric
To comprehensively evaluate our method, we conduct experiments using both synthetic data and real-world data. The synthetic point clouds we used are derived from ShapeNet [70]. We render a single depth map for each object and then back-project it into 3D space to obtain the partial point cloud. The complete point clouds for evaluation are uniformly sampled from the objects’ surfaces. Following PCN [15], we select 30974 objects across 8 categories: airplane, cabinet, car, chair, lamp, couch, table, and watercraft. For each category, we employ 100/150 objects for validation/testing, and the rest is used for training. We set the point numbers of partial input and complete shape to 2048 and 16384, respectively. The evaluation metric for synthetic data is CD, which is widely used in point cloud completion tasks.
@@ IV-B Implementation Details
The point numbers of $\mathbf{P}_{in}$, $\mathbf{P}_{c}$, and $\mathbf{P}_{out}$ are 2048, 2048 and 16384, respectively. We employ neighbor numbers of 16 and 24 for calculating position and curvature encodings, respectively. Initial point radii for the DARE algorithm are set to 0.008 for the plane category and 0.03 for other categories. The loss weights are all set to 1. The initial learning rate is 0.0001, and it decays by 0.5 for every 200 epochs. The training process takes 600 epochs to converge. The experiments are conducted on an NVIDIA RTX 3090 GPU with Intel(R) Xeon(R) CPUs.
@@ IV-C Comparison with State-of-the-Art Methods
IV-C1 Results on synthetic data
We first conduct experiments on the synthetic dataset and compare the results with other state-of-the-art methods, including fully supervised methods (i.e., FoldingNet [40], PCN [15], TopNet [41], AnchorFormer [17]), unpaired methods (i.e., PCL2PCL [25], KTNet [29], USSPA [30]) and self-supervised methods (i.e., ACL-SPC [64], P2C [65]). The quantitative results presented in Table I suggest that our self-supervised method achieves an impressive average CD of 9.08, significantly outperforming other self-supervised methods and two unpaired approaches (i.e., PCL2PCL and USSPA) by a large margin, and even demonstrating competitive results compared to some fully supervised methods. Regarding the category-specific results, our method significantly outperforms two self-supervised methods across five categories. Specifically, the CD value of MAL-SPC under the lamp category surpasses the results of P2C and ACL-SPC by nearly 5 and 20, respectively. Besides, the CD values of MAL-SPC are lower than those of self-supervised benchmarks by 2-4 under cabinet, couch, and table categories. Furthermore, even with the prior knowledge of complete 3D point clouds, the results of the two unpaired methods, PCL2PCL and USSPA, lag significantly behind our method across multiple categories. In addition, compared to the supervised method FoldingNet, our model achieves a CD difference of less than 3. These results demonstrate the strong performance and effectiveness of our method.
@@ IV-C1 Results on synthetic data
TABLE III: Comparison of FLOPs and model size.
@@
IV-C2 Results on real-world scans
In order to validate the robustness and effectiveness of our method in real-world environments, we evaluate its performance on two challenging datasets: ScanNet and KITTI. As demonstrated in Table II, compared to ACL-SPC and P2C, our MAL-SPC yields the best UCD on all three categories and the best UHD on the car and table categories. Our method only slightly falls behind the second-best approach in terms of UCD on the chair category. Moreover, regarding the results on the car category, MAL-SPC even surpasses the unpaired and supervised methods by a large margin. Figure 5 shows the quantitative results of our model and other state-of-the-art benchmarks. Similar to its superior performance on synthetic data, our method predicts more reasonable shapes and better preserves the original inputs, while generating significantly fewer noise points.
TABLE IV: Quantitative comparison of the effectiveness of each module in our framework. “✓” and “✕” represent the model with or without the component, respectively. “PE” and “CE” denote the position encoding and the curvature encoding, respectively. “MAN” means the multi-view adversarial network, and “DARE” indicates the density-aware radius estimation algorithm. The bold numbers represent the best results.

### v2

@@ III-B Multi-view Adversarial Learning
Objects of the same category share many similar structures, and therefore, their shape distributions are close. To make the generated complete point cloud as realistic and reasonable as possible, we develop an adversarial learning module to learn the category-specific shape knowledge from the real depth maps. Specifically, we first employ a silhouette renderer to render $\mathbf{P}_{out}$ and $\mathbf{P}_{c}$ into two silhouette maps $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ from the same viewpoint as $\mathbf{D}_{0}$. $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ are supervised by $\mathbf{D}_{0}$. Then we leverage a depth renderer to render $\mathbf{P}_{out}$ into depth maps from different viewpoints. These depth maps and those from the category-specific depth image bank are fed to a discriminator, which consists of five 2D convolutional layers, to discriminate whether they are generated or real samples. It is important to note that the depth image bank is composed of single-view maps, each corresponding one-to-one with the input partial shape, without including samples from additional viewpoints.
@@
Our differentiable rendering pipeline is derived from [70]. Given the intrinsic and extrinsic camera parameters, [70] first transform the point cloud from world to camera coordinates and then project it into the 2D image plane. During rasterization, each point is rendered as a circular splat with a predefined radius in screen space, where the opacity of the splat decreases away from its center. The value of each pixel in the silhouette maps is computed by blending the information from the K nearest points along the z-axis whose splat regions overlap that pixel. For depth maps, each pixel’s value is the smallest z-coordinate of the K nearest points.
@@
However, [70] assigns a fixed and identical radius to each point splat. In fact, the density of 2D points projected by a point cloud with a fixed point number from different viewpoints is different. For example, as shown in Figure 3, when there is a large foreground area of a rendered image, the density of the projected points is small (we call these low-density depth maps). In such cases, a small radius cannot cover all areas of the visible surface and thus many “hole pixels” will appear in the rendered depth map. These “hole pixels” are inaccurately represented by the depth values of the points that should have been occluded. In contrast, in scenarios with few foreground pixels, the density of points is large (we call these high-density depth maps). In this case, a small radius suffices to cover the entire visible surface, and thus, the rendered depth map is more realistic and smooth. Therefore, a fixed point radius is inadequate for rendering under different viewpoints. Intuitively, using a larger point radius might produce smoother depth maps in both high point density and low point density scenarios. However, in this way, the projected points cover a larger area on the 2D plane, potentially extending the edges of the foreground pixels beyond the original boundaries of the object.
@@
To this end, we introduce a density-aware radius estimation (DARE) algorithm to adaptively determine a suitable radius for points from each viewpoint. First, we predefine an initial radius for the generated points and then render them into a coarse depth map. Note that the predefined point radius should cover the projected area. We then calculate the number of foreground pixels on this depth map, denoted as $A$, and subsequently, the density of the projected points on the depth map is defined as
@@
where $r$ denotes the point radius. Note that the parameter $\eta$ can be pre-estimated based on the initial point radii. Specifically, for each category, the density of projected points could be approximated using the average foreground pixel number of all the depth maps in the dataset, as specified in Formula 6. Following this, $\eta$ can be computed according to Formula 7. In this way, the radius of the point cloud dynamically adjusts according to the density of projected points, yielding a smoother depth map.
@@ III-C Loss Functions
Furthermore, we use a rendering loss to supervise the predicted point cloud in 2D space. First, we binarize the depth image $\mathbf{D}_{0}$ to obtain the 2D silhouette ground truth $\mathbf{S}_{0}$ from the input viewpoint. Subsequently, we compute the Mean Squared Error (MSE) loss between $\mathbf{S}_{0}$ and $\mathbf{S}_{out}$. Note that due to the sparsity of the coarse point cloud $\mathbf{P}_{c}$, the foreground pixels of $\mathbf{S}_{c}$ are also sparse and discrete, whereas the foreground of $\mathbf{S}_{0}$ is continuous and dense. To solve this problem, we propose a Masked MSE loss, which only calculates the MSE loss on the foreground pixels of the rendered image. This trick confines the projected points of the rendered image to be accurately positioned.
@@
[…]
Methods Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
ACL-SPC[66] 13.43/1.81 2.75/0.63 19.66/2.06 4.05/1.82 8.88/2.49 33.14/1.59 10.00/3.03 18.44/1.79 10.59/1.14
P2C [67] 9.28/3.00 1.33/0.52 10.02/2.71 2.97/1.96 7.82/5.17 18.10/1.33 17.05/7.77 11.97/2.98 5.01/1.56
Ours 6.33/2.74 1.43/0.92 8.08/2.56 4.28/1.63 6.27/3.87 10.52/3.99 7.71/2.44 7.23/4.42 5.17/2.10
IV Experiments
IV-A Dataset and Evaluation Metric
To comprehensively evaluate our method, we conduct experiments using both synthetic data and real-world data. Since existing datasets lack the initial depth maps for partial point clouds, we generate a new synthetic dataset from ShapeNet [72]. Specifically, we render a single depth map for each object and then back-project it into 3D space to obtain the partial point cloud. To better align with human perspectives, we set the elevation angle for scanning shapes between -30 and 30 degrees. The complete point clouds for evaluation are uniformly sampled from the objects’ surfaces. Following PCN [15], we select 30974 objects across 8 categories: airplane, cabinet, car, chair, lamp, couch, table, and watercraft. For each category, we employ 100/150 objects for validation/testing, and the rest is used for training. We set the point numbers of partial input and complete shape to 2048 and 16384, respectively. Besides, we evaluate our model on widely used benchmarks, including PCN [15], 3D-EPN [34] and CRN [24]. The evaluation metric for synthetic data is CD. Additionally, following previous self-supervised approaches [65, 66], we report each component of CD, i.e., Precision and Coverage.
[…]
Dataset Methods Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
PCN ACL-SPC[66] 15.81 4.29 22.76 5.80 12.06 34.21 14.13 21.52 11.75
P2C [67] 17.14 2.18 20.64 4.77 13.76 19.06 55.18 14.54 7.00
Ours 9.77 3.38 11.20 6.32 10.20 14.52 14.12 10.71 7.72
3D-EPN ACL-SPC[66] 20.75 4.57 29.34 9.00 17.83 43.79 17.28 30.61 13.63
P2C [67] 14.54 2.58 19.35 8.00 13.39 22.67 22.11 19.76 8.51
Ours 18.69 3.26 30.23 10.35 24.00 24.76 18.18 28.90 9.86
CRN ACL-SPC[66] 16.81 4.56 24.30 7.03 13.11 34.71 15.53 22.86 12.40
P2C [67] 18.18 2.47 22.54 5.91 14.88 19.70 56.85 15.48 7.63
Ours 10.72 3.70 12.80 7.42 11.10 15.20 15.45 11.72 8.37
@@ IV-B Implementation Details
The point numbers of $\mathbf{P}_{in}$, $\mathbf{P}_{c}$, and $\mathbf{P}_{out}$ are 2048, 2048 and 16384, respectively. We employ 16 and 24 neighbors for calculating position and curvature encodings, respectively. Initial point radii for the DARE algorithm are set to 0.008 for the plane category and 0.03 for other categories. The loss weights are all set to 1. The initial learning rate is 0.0001, and it decays by 0.5 for every 200 epochs. The training process takes 600 epochs to converge. The experiments are conducted on an NVIDIA RTX 3090 GPU with Intel(R) Xeon(R) CPUs.
@@ IV-C Comparison with State-of-the-Art Methods
IV-C1 Results on synthetic data
We first conduct experiments on the synthetic dataset and compare the results with other state-of-the-art methods, including fully supervised methods (i.e., FoldingNet [41], PCN [15], TopNet [42], AnchorFormer [17]), unpaired methods (i.e., PCL2PCL [25], KTNet [29], LTUPCN [63], USSPA [30]) and self-supervised methods (i.e., ACL-SPC [66], P2C [67]). We train our model in two modes: one is training separately for each category, and the other is training all eight categories together while the adversarial depth maps for each object are from the same category. In contrast, all compared methods are only trained for each category. The quantitative results presented in Table I suggest that our self-supervised method achieves an impressive average CD of 9.08, which outperforms other self-supervised methods and two unpaired approaches (i.e., PCL2PCL and USSPA) by a large margin, and even demonstrating competitive results compared to some fully supervised methods. Regarding the category-specific results, our method significantly outperforms two self-supervised methods across five categories. Specifically, the CD value of MAL-SPC under the lamp category surpasses the results of P2C and ACL-SPC by nearly 5 and 20, respectively. Besides, the CD values of MAL-SPC are lower than those of self-supervised benchmarks by 2-4 under cabinet, couch, and table categories. Furthermore, even with the prior knowledge of complete 3D point clouds, the results of the two unpaired methods, PCL2PCL and USSPA, lag behind our method across multiple categories. In addition, compared to the supervised method FoldingNet, our model achieves a CD difference of less than 3.
As shown in the last two rows of Table I, the results of our method trained separately for each category are generally better than training them together. However, the latter also outperforms other self-supervised methods. Table II shows that our method achieves the best Precision and Coverage across most categories, and outperforms other self-supervised methods in terms of average Precision. Although it falls slightly behind ACL-SPC in Coverage, the difference remains minimal. These results indicate that our method predicts highly complete point clouds while effectively matching local shapes. Table III indicates that our method leads on the PCN and CRN results, with the average CD surpassing the second place by 6.04 and 6.09, respectively. However, the performance of our method and ACL-SPC shows a significant decline on the 3D-EPN dataset. To simulate the human perspective, we set the elevation angle for scanning objects between -30 and 30 degrees when generating our dataset, whereas the partial point clouds in 3D-EPN are scanned from arbitrary angles. This difference in viewpoints introduces a domain gap between the datasets, resulting in the underperformance of our method and ACL-SPC. Note that the performance of P2C increases on the 3D-EPN dataset, indicating that it is more robust to viewpoint variations of partial shape scan.
Moreover, we evaluate the performance of our method in the unpaired case. Specifically, we replace our MAN module with the MLP-based discriminator from USSPA and use the predicted point cloud and an unpaired complete shape as discrimination targets. As shown in Table I, our method achieved favorable results in the unpaired case, with the average CD significantly outperforming PCL2PCL and USSPA. However, the results slightly lag behind our self-supervised approach. We believe that, on the one hand, our PRN effectively learns the common 3D structural features within each object and category, thereby reducing the impact of the unpaired complete point cloud. On the other hand, unlike unpaired methods, our MAN leverages a 2D network as the discriminator. It is well-known that 2D neural networks generally possess stronger learning capabilities compared to 3D MLPs, and our method is able to effectively utilize this advantage.
@@ IV-C1 Results on synthetic data
[…]
Methods Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
ACL-SPC [66] 25.04 10.47 25.63 11.33 25.35 61.68 26.93 23.71 15.29
P2C [67] 18.67 5.62 16.05 16.00 15.68 36.29 25.81 19.06 14.83
Ours 16.93 4.34 12.82 8.07 19.20 44.09 15.85 18.84 12.28
[…]
Methods Plane Cabinet Car Chair Lamp Couch Table Watercraft
ACL-SPC [66] 0.74/10.84 17.16/44.00 6.40/8.99 7.46/14.71 12.23/50.15 9.08/37.38 14.45/60.62 6.48/23.54
P2C [67] 0.53/4.91 8.75/16.58 4.01/7.54 5.60/35.13 6.16/21.83 9.09/104.86 3.98/57.88 3.67/13.15
Ours 0.87/5.97 11.17/12.88 6.78/14.39 5.39/16.82 9.36/13.82 9.22/11.46 7.67/41.41 3.57/11.23
IV-C2 Results on real-world scans
In order to validate the robustness and effectiveness of our method in real-world environments, we evaluate its performance on two challenging datasets: ScanNet and KITTI. As demonstrated in Table IV, compared to ACL-SPC and P2C, our MAL-SPC yields the best UCD on all three categories and the best UHD on the car and table categories. Our method only slightly falls behind the second-best approach in terms of UCD on the chair category. Moreover, regarding the results on the car category, MAL-SPC even surpasses the unpaired and supervised methods by a large margin. Figure 5 shows the quantitative results of our model and other state-of-the-art benchmarks. Similar to its superior performance on synthetic data, our method predicts more reasonable shapes and better preserves the original inputs, while generating significantly fewer noise points.
Figure 6: Comparison of visual results on the severely occluded dataset. (a) Input partial points. (b) P2C [67]. (c) ACL-SPC [66]. (d) Ours. (e) Ground truth.
IV-C3 Results on severely occluded objects
To assess our model’s ability to complete objects with severe occlusions, following [16], we remove half of the points that are furthest from the camera in the partial scans and then use these data to train and test our model. As indicated in Table V, the performance of the three self-supervised methods significantly decreases when completing objects with severe occlusions. However, our method still leads the others, achieving the best results in terms of average CD and the CD across six categories. The visual results in Figure 6 demonstrate that our method can still infer the complete shape under extreme occlusion conditions. For example, even when only parts of an airplane’s wing and tail are visible, our method can still reconstruct the fuselage. We believe this is because our model learns category-specific geometric patterns to complete missing parts.
IV-C4 Results on long-tail objects
Our model takes advantage of the common geometries within each category. To test its ability to reconstruct shapes with rare structures, we select long-tail objects of each category for experiments. Specifically, we use a pre-trained point cloud reconstruction model to generate a global feature for each object, and then calculate the cosine similarity between each global shape and others within the same category. The ten samples with the lowest average similarity are selected as the long-tail set for that category, while the ten with the highest average similarity form the common set, representing objects with the most typical geometric structures. As Table VI demonstrates, the performance of the three self-supervised methods drops significantly on the long-tail set, as they rely on common geometric patterns to infer missing structures without supervision from complete point clouds. In addition, our method demonstrates stable performance on the cabinet and couch categories, while on the table category, the performance difference between the two sets is the largest, reaching nearly 34.
TABLE VII: Quantitative results on unseen categories. “Major Categ.” and “Minor Categ.” represent the two categories in ShapeNet-34 with the largest and smallest number of objects, respectively. The bold numbers represent the best results.
Methods Average Major Categ. Minor Categ.
bowl washer cap keyboard
ACL-SPC [66] 230.54 59.79 431.94 132.45 240.59
P2C [67] 119.96 52.81 249.57 83.03 20.66
Ours 98.89 49.84 196.15 83.84 13.02
IV-C5 Results on novel categories
To evaluate the generalization ability of our model, we train it on all eight categories and then test it on novel categories. Specifically, the testing data are derived from ShapeNet-34 [16], which comprises 34 categories from ShapeNet. As shown in Table VII, the self-supervised methods generally perform poorly on the unseen dataset, indicating their limitations in generalization. This is because, without complete supervision, the self-supervised approaches can only capture some typical geometric structures within the training categories. Even so, our method outperforms ACL-SPC and P2C and demonstrates particularly good performance on the ’keyboard’ category.
TABLE VIII: Comparison of FLOPs and model size.
@@ IV-C5 Results on novel categories
IV-C6 Comparison of computational complexity and model size
Table VIII provides a comprehensive comparison of the computational complexity, measured in Floating Point Operations (FLOPs), and the model size for our proposed method against several state-of-the-art approaches: USSPA, ACL-SPC, and P2C. Our model demonstrates a well-balanced approach in terms of both computational complexity and model size. With 15.10G FLOPs, it is considerably more efficient than USSPA. In terms of model size, our method, with 94.79M parameters, remains manageable and smaller than both ACL-SPC and P2C. This balance makes our model an attractive option for practical applications where both computational resources and memory capacity are constrained.
Figure 7: Examples of the failure cases. (a) Partial inputs. (b) Ours. (c) Ground truth.
IV-C7 Failure cases
In Figure 7, we present some examples of failure cases. When the shape of the target object is rare within its category or the occlusion is severe, our method struggles to reconstruct these unique structures, instead erroneously generating more common structures. This is because, without supervision, our method heavily relies on learning the common geometric features within the category to reconstruct missing structures, which is a constraint shared by current self-supervised methods.
IV-D Ablation Study
IV-D1 Effects of each module
TABLE IX: Quantitative comparison of the effectiveness of each module in our framework. “✓” and “✕” represent the model with or without the component, respectively. “PE” and “CE” denote the position encoding and the curvature encoding, respectively. “MAN” means the multi-view adversarial network, and “DARE” represents the density-aware radius estimation algorithm. The bold numbers represent the best results.